OpenStudy (anonymous):

Find an equation of the line with given slope that passes through the given point. write the equation in ax+by=c. m=8/5, (9,-2)

OpenStudy (anonymous):

first use the slope intercept form first: y=mx + c and plug in the values to solve for c

OpenStudy (texaschic101):

do you understand what jayz657 said ?

OpenStudy (texaschic101):

let us know when you are ready to participate

OpenStudy (anonymous):

not really

OpenStudy (texaschic101):

jayz657...I do not want to take over your question unless you want me to. Do you want to take this, or do you want me to ?

OpenStudy (anonymous):

i'll try to explain so slope intercept form is y = mx + c where y is the y value of the coordinate m is the slope, in this case it is 8/5 and x is the x value coordinate and remember points are written in this form (x,y) all you need to do is substitute everything into the slope intercept form and try to solve for the missing variable, which is c.

OpenStudy (anonymous):

my answer choices are 8x+5y=-82 5x-8y=-82, 8x-5y=82 or 8x+5y=82 so i dont understand how what oyu are saying will work

OpenStudy (texaschic101):

sub in your info into the formula...y = mx + c m = 8/5 (9,-2) x = 9 and y = -2 now what can you do with that Ariel ?

OpenStudy (texaschic101):

y = mx + c -2 = 8/5(9) + c do you understand how I got that ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

can you solve for c ?

OpenStudy (anonymous):

no it gives no solution

OpenStudy (texaschic101):

you have to find the solution... -2 = 8/5(9) + c -2 = 72/5 + c (subtract 72/5 from both sides) -2 - 72/5 = 72/5 - 72/5 + c (combine like terms) -2 - 72/5 = c -10/5 - 72/5 = c -82/5 = c do you understand so far ?

OpenStudy (anonymous):

yeah

OpenStudy (texaschic101):

now your equation is : y = 8/5x - 82/5 still with me ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

now we need the standard form Ax + By = C y = 8/5x - 82/5 (you need to get rid of the fractions. so multiply the entire equation by the LCD which is 5) giving you : 5y = 8x - 82 (subtract 8x from both sides) -8x + 5y = -82 (multiply by -1 to get x positive) 8x - 5y = 82 <-- standard form any questions ?

OpenStudy (anonymous):

no thank you so much

OpenStudy (texaschic101):

glad to help. I really hope you understand this because you will probably get more problems like this. I hope you just didn't want the answer

OpenStudy (anonymous):

no im glad that it saves our old problems so that i can come back to this when i do get another problem that is similar